Water vapor accounts for part of the gas volume and S8 lle volum u sued must be corrected for the water vapor present. The vawhere we assume that the pressure exerted by the gas sample, Pgas, is atmospheric pressure. Thus, because pressure and volumePiatal, mm Hg : 745.490 T°C Vgas, mL PH20, torr Pair/ torr Vair, mL 66.8 7.95 205.119 64.6 7.59 179.452 62.4 7.28 163.896 60.Find P air and V air ..
Answer #1

T°C VGAS, Puzo, Vair, mL Pa torr Par = P -PH2O mL torr V PE-PH20)/PEX Vair, ml (result of previous column) 5.7625 66.8 7.95 2

Y-Values y= 0.015x + 4.760 R = 0.985 volume of air Y-Values - Linear (Y-Values) -300 100 -200 -1000 temperature in degree cel

y= 0.015x + 4.760 R= 0.985 volume of gas in ml niiniiniiniin + 10 50 60 70 80 20 30 40 temperature in degree celsius

y = 0.015x + 4.760

put y = 0 to find x intercept

0 = 0.015x + 4.760

X = -4.760/0.015 = -317.340C

So value of absolute zero according to best fit line is -317.340C

Actual value absolute zero is -273.150C